AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 06:33:40 2025
Author
*
:
Subject
*
:
<p>07:10</p> <p>FEE temperatures OK</p> <p>System checks OK<br /> Merger and tape server OK<br /> Statistics look OK,<br /> Rates in the detectors from the mointoring code indicate we have beam on target.</p> <p>Detector leakge current appears to be following same pattern as yesterday -> temperature (attachment 1)</p> <p>Merger stats (attachment 2) - quite a few more pause / resume statements than normal - noise last night?</p> <p>Target density appears stable over the last 3 hours ~3.8e11 at/cm2, ring pressure stable YR09 ~1.5e-10 mbar</p> <p>Beam injection looks stable over the last few hours.</p> <p>09:08</p> <p>FEE temperatures OK -aida03 and aida09 0.5 C over 65C warning.</p> <p>system checks OK</p> <p>ADC data item stats OK</p> <p>Spectrum browser rate (attachment 3)</p> <p>Merger and Tape server OK</p> <p>Monitoring xy and energy histograms since zero'd last night (attachments 4+5)</p> <p>Target density slight drop to ~3.5e11 at/cm2. Pressure in the ring stable. CB attempting to optimise target density and ring pressure.</p> <p>Beam injection into the ring still good</p> <p>11:04</p> <p>FEE temperatures OK</p> <p>System checks OK</p> <p>ADC data items and spectrum browser rates OK</p> <p>Target density ~4.6e11 at/cm2, ring pressure looks OK ~1e-10 mbar</p> <p>Beam injection still good.</p> <p>Merger and tape server OK</p> <p>12:45</p> <p>Beam injection over last 2 hours has been stable</p> <p>Target density 4.75e11 and stable. ring pressures are stable.</p> <p>FEE temperatures OK</p> <p>System checks OK</p> <p>ADC data items stats and spectrum browser rates look OK</p> <p>Merger and tape server OK</p> <p>14:01</p> <p>We wish to test the beam ahlo on injection so are ending run</p> <p>Run stopped.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown